# DualMAR: Medical-Augmented Representation from Dual-Expertise Perspectives

### v1

@@ Abstract
Electronic Health Records (EHR) has revolutionized healthcare data management and prediction in the field of AI and machine learning. Accurate predictions of diagnosis and medications significantly mitigate health risks and provide guidance for preventive care. However, EHR driven models often have limited scope on understanding medical-domain knowledge and mostly rely on simple-and-sole ontologies. In addition, due to the missing features and incomplete disease coverage of EHR, most studies only focus on basic analysis on conditions and medication. We propose DualMAR, a framework that enhances EHR prediction tasks through both individual observation data and public knowledge bases. First, we construct a bi-hierarchical Diagnosis Knowledge Graph (KG) using verified public clinical ontologies and augment this KG via Large Language Models (LLMs); Second, we design a new proxy-task learning on lab results in EHR for pretraining, which further enhance KG representation and patient embeddings. By retrieving radial and angular coordinates upon polar space, DualMAR enables accurate predictions based on rich hierarchical and semantic embeddings from KG. Experiments also demonstrate that DualMAR outperforms state-of-the-art models, validating its effectiveness in EHR prediction and KG integration in medical domains.
Keywords EHR Prediction  $\cdot$ Diagnosis KG  $\cdot$ Medical-domain knowledge  $\cdot$ LLM  $\cdot$ Pretraining  $\cdot$ Lab results
@@ 1 Introduction
The digitization of patient information such as EHR has transformed healthcare in terms of data storage, information retrieval, and computational pattern recognition. EHR-based prediction models, including those for diagnosis, heart failure, and readmission prediction tasks (Shickel et al., 2017), leverage EHR data to forecast patient outcomes. In addition to improving predictive accuracy, a major focus of current research is to build models that understand medical domains from both individual health records (Poulain and Beheshti, 2024) and public knowledge bases (Jiang et al., 2023).
Individual health records preserve hidden knowledge of medical concepts, such as disease complications. Many models use EHR data to uncover disease progression patterns and learn medical embeddings, which serve as local-domain priors for downstream tasks. Meanwhile, transformer-based models following G-BERT (Shang et al., 2019) rely on self-supervised tasks to fully utilize EHR data, including single admissions, enhancing training efficiency and robustness of models. However, these methods typically focus on condition, treatment, and medication concepts, often overlooking critical information like lab tests.
(a) Generating Augmented Bi-hierarchical KG.
(b) The Dual-Expertise Module for EHR Prediction.
Figure 1: Overview of the proposed DualMAR framework.
From another perspective, public medical knowledge have been used in machine learning to enhance understanding of medical concepts Choi et al. (2017); Ma et al. (2018). Recent studies still primarily focused on simple hierarchical ontologies, like ICD-9 hierarchy Lu et al. (2023), showing that even simple and sole relations can help models get state-of-the-art performance in predictive healthcare. However, KGs with more complex and diverse relations remain under-explored in EHR models. To the best of our knowledge, GraphCare Jiang et al. (2023) is one of the preliminary attempts to use large language models such as GPT-4 Achiam et al. (2023) to generate KGs for EHR predictions, outperforming most existing models. Experimental results demonstrate that KGs refine prediction through comprehensive understanding for medical concepts, making them a promising direction for augmenting EHR models, though some unsolved challenges remain:
@@
    Data Quality of KG: GraphCare suffers from data quality issues with its generated triples. First, triples generated by GPT in GraphCare can lead to hallucinations upon zero-shot learning. Additionally, the final KG used in GraphCare is constrained by its reliance on the CCS Wei et al. (2017) encoding system, limiting its scope (# diseases: 285) for more specific diagnostic tasks.
@@
    Integration of KG into EHR Models: First, GraphCare initializes each patient with a personalized KG for prediction, which imposes demands on GPU memory and limits scalability as KG sizes increase. Second, GraphCare primary focus on the semantic information of medical concepts leads to a partial understanding within the KG, which, combined with its lack of hierarchical features, can result in biased predictions for EHR tasks.
To address these challenges in integrating KGs into predictive models, we propose leveraging both medical domain knowledge from generated KGs and a novel pretraining proxy task in EHR, which enables models to take advantage of both “knowledge scholar” and “local expert”. Our proposed method, DualMAR, is a framework which not only enhance embedding by retrieving both contextual and hierarchical information from KGs but also further refine medical representation through EHR graphs and pretraining modules. The main contributions of this work are as follows:
  • “Knowledge Scholar”: We establish a diagnosis-driven KG across public medical databases and design a bi-hierarchical structure that organizes core domains based on phenotype and condition codes, further enriching it with verified textual information through in-context learning with LLMs. We then employ polar-space projection to retrieve embeddings that capture both semantic and hierarchical information from generated KG.
  • “Local Expert”: We employ a Graph Learning module using a disease-lab EHR graph to enhance embeddings of medical codes, and introduce a novel proxy task to augment patient hidden representations from a laboratory perspective, fully leveraging the entire EHR dataset.
  • “Dual-expertise Perspective”: An encoder-decoder architecture is adopted for integrating KG into the predictive model. Embeddings provided by the “Knowledge Scholar” serve as prior knowledge for the “Local Expert” who continually refines these representations. The “Local Expert” converts patient admission history into laboratory embeddings through the encoder, resulting in accurate predictions across various downstream tasks.
@@ 2 Related Work
Deep learning models have been extensively applied in EHR prediction to provide guidance for preventive care. Both RNN-based Choi et al. (2016a, b); Ma et al. (2017); Bai et al. (2018); Ma et al. (2020a) and CNN-based models Nguyen et al. (2017); Ma et al. (2020b) have been developed to capture temporal and spatial dependencies for health risk predictions. With the advent of graph neural networks (GNNs), GRAM Choi et al. (2017) and its successors Liu et al. (2020); Wu et al. (2021); Lu et al. (2023) incorporate graph learning to further enhance hidden representation for medical concepts. Meanwhile, previous work such as G-BERT Shang et al. (2019) takes advantage of transformers and their variants to fully utilize EHR data (i.e. single-admission patients) through various proxy tasks Prakash et al. (2021); Rasmy et al. (2021); Poulain and Beheshti (2024).
@@
Predictive models that leverage graph and transformer architectures have achieved state-of-the-art performance in various EHR tasks, highlighting their potential to advance predictive healthcare with existing knowledge. This knowledge can be divided into local individual health data and public knowledge bases, as outlined below:
  • Individual Health Data: Prior work such as CGL Lu et al. (2021) has constructed EHR graphs from EHR data based on co-occurrence of medical concepts in admission records to capture hidden relations among diseases and improve health outcome predictions. Encoder-decoder models following G-BERT Shang et al. (2019), focus on customized pretraining tasks to enhance the understanding of medical patterns in EHR data. While individual knowledge aids in adapting models to different datasets, it relies on the size of samples and the quality of data.
@@
  • Public Knowledge Bases: External knowledge is derived from public data sources without individual information. Researchers have explored simple ontologies like the ICD-9 hierarchy, Retrieval-Augmented methods like RAM-EHR Xu et al. (2024) for unstructured medical descriptions, and KGs for predictive models like GraphCare Jiang et al. (2023). However, these approaches often focus solely on either hierarchical or semantic information, leading to a narrow understanding of medical concepts and potential information loss, and how to effectively combine both information in models is necessary to enhance robustness of predictive healthcare.
@@
In addition to leveraging medical-domain knowledge, incorporating both structured and unstructured features can enhance predictions. MiME Choi et al. (2018) and GCT Choi et al. (2020) use graph structures with lab results to augment representations, while CGL Lu et al. (2021) and MedGTX Park et al. (2022) integrate unstructured clinical notes to highlight the value of additional information. However, most approaches rely heavily on these multi-aspect features, limiting their applicability to data without those features. Our proposed framework integrates laboratory features as auxiliary data in graph construction and proxy tasks, enhancing predictions even when no desired feature is available.
@@ 3 Methodology
We present DualMAR, a framework designed for predictive healthcare that aims to achieve a comprehensive understanding of diagnosis-related entities from both external (public) and local (individual) knowledge bases. An overview of the proposed framework is shown in Figure 1(b).
@@ 3.1 General Notation
An EHR dataset $\mathcal{S}$ is a collection of temporal admission records of $N$ patients $\{\mathbf{p}_{1},\mathbf{p}_{2},...,\mathbf{p}_{N}\}$, and $\mathbf{p}_{u}$ is denoted as certain patient $u$. Such patient can be also represented as a sequence of $T_{u}$ admission records $(V_{1}^{u},V_{2}^{u},..,V_{T}^{u})\in\mathbf{p}_{u}$ in chronological order. We omit $u$ to identify different patients in following sections and explain our framework using single patient as example to avoid misunderstanding. We denote the entire set of medical concepts as $\mathcal{C}=\{c_{1},c_{2},...,c_{|\mathcal{C}|}\}$ where $|\mathcal{C}|$ is the vocabulary size of medical concepts. Each admission, like $t$-th admission $V_{t}\in(V_{1},V_{2},..,V_{T})$, contains a subset of $\mathcal{C}$. Note that, We consider diseases and lab results as medical concepts in proposed framework, and we denote diseases $\{d_{1},d_{2},...,d_{|\mathcal{D}|}\}$ and lab results $\{l_{1},l_{2},...,l_{|\mathcal{L}|}\}$ within vocabulary $\mathcal{D},\mathcal{L}\subset\mathcal{C}$.
3.2 Generate “Knowledge Scholar”
In general, the proposed model integrates external knowledge from professional medical datasets and LLM-retrieved KGs based on verified clinical text. Figure 1(a) illustrates how to augment structured KG with textual information.
(1) HKG Construction: Healthcare KGs (HKGs) are healthcare-specific KGs that span multiple datasets (Cui et al., 2023), offering a more comprehensive understanding than all-in-one ontologies like UMLS Bodenreider (2004). We focus on constructing HKGs that link diseases, drugs, and phenotypes to enable precise clinical predictions. The Human Phenotype Ontology (HPO) and the ICD-9 hierarchy serve as the foundation for our KG, with bi-hierarchical structure enhancing its robustness. Inspired by previous healthcare KGs (Su et al., 2023; Chandak et al., 2023; Ghorbanali et al., 2023), we also incorporate updated and verified sources like Drug Central, DrugBank, and SIDER to create a highly connected HKG. We denote the resulting healthcare KG, sourced from existing databases, as $\mathcal{G}_{M}$, containing $M$ tuples of head, relation, and tail. More details can be found in Appendix A.
(2) LLM-KG Generation: We extract textual features for specific diseases from widely-used knowledge bases and then employ a template with instruction, example, and prompt to harness LLMs’ in-context learning capabilities. For instance, when generating a KG for a diagnostic concept, the LLM generates triples like “[Heart Failure, CAUSED_BY, Hypertension]” based on the collected text. Iterative prompting and re-reading strategies are used to refine outputs. The aggregated results form an LLM-generated KG $\mathcal{G}_{N}$. Further details on knowledge sources and prompting are in Appendices A.2 and A.3 respectively.

### v2

# Bridging Stepwise Lab-Informed Pretraining and Knowledge-Guided Learning for Diagnostic Reasoning

@@ Abstract
Despite the growing use of Electronic Health Records (EHR) for AI-assisted diagnosis prediction, most data-driven models struggle to incorporate clinically meaningful medical knowledge. They often rely on limited ontologies, lacking structured reasoning capabilities and comprehensive coverage. This raises an important research question: Will medical knowledge improve predictive models to support stepwise clinical reasoning as performed by human doctors? To address this problem, we propose DuaLK, a dual-expertise framework that combines two complementary sources of information. For external knowledge, we construct a Diagnosis Knowledge Graph (KG) that encodes both hierarchical and semantic relations enriched by large language models (LLM). To align with patient data, we further introduce a lab-informed proxy task that guides the model to follow a clinically consistent, stepwise reasoning process based on lab test signals. Experimental results on two public EHR datasets demonstrate that DuaLK consistently outperforms existing baselines across four clinical prediction tasks. These findings highlight the potential of combining structured medical knowledge with individual-level clinical signals to achieve more accurate and interpretable diagnostic predictions. The source code is publicly available on https://github.com/humphreyhuu/DuaLK.
{IEEEkeywords}
Diagnosis Prediction, Diagnosis Knowledge Graph, LLM, Clinical Reasoning, Stepwise, Lab Tests
@@ 1 Introduction
The digitization of patient information such as EHR has transformed healthcare in terms of data storage, information retrieval, and computational pattern recognition. EHR-based prediction models, including those for diagnosis [1], heart failure [2], and readmission prediction tasks [3], leverage EHR data to forecast patient outcomes. To reduce bias risks of predictions based on data patterns rather than clinical understanding, it is essential to integrate medical domain knowledge from both individual health records and public medical knowledge into the predictive process. (1) Individual knowledge can help model capture progression patterns (e.g. disease complications) of personal health status on EHR datasets [4, 5, 6]; (2) Public knowledge bases and clinical websites have been used to enhance the understanding of medical concepts [7, 8, 9], which reveals that even simple hierarchical structures of diseases can improve prediction performance.
Despite recent progress, performing diagnostic decision-making in a manner similar to physicians remains a challenging goal, with several challenges ahead:
@@
    Underutilized public medical knowledge with limited quality. While external knowledge beyond the training data has been widely utilized by previous approaches, these models often rely on simple ontologies (e.g., the hierarchy in ICD-9-CM [10]), which fail to capture the comprehensive medical knowledge required for clinical prediction [7, 4]. Textual and other unstructured features [5, 11, 12] have also been explored to improve predictive performance, but semantic ambiguity and irrelevant context remain challenges.
@@
    Lack of diagnostic reasoning when using EHR data. Some studies [4, 13, 14] take advantage of customized pretraining tasks to fully exploit EHR data, overlooking the causal and procedural nature of clinical decision-making. There are also some models using co-occurrence graphs to learn medical interactions (e.g., complications) [1, 5, 6], but the relations they capture are simple associations, failing to reflect the multi-step diagnostic reasoning that physicians typically follow.
  3. 3.
    Inefficiency of integrating medical knowledge in EHR. Most models rely on either public knowledge or individual health data independently, as combining both may incur high computational costs during deployment [15]. However, models that rely on a single source of data fail to capture generalized patterns, leading to biased predictions and degraded robustness.
To overcome these limitations, a stepwise and knowledge-informed framework is required to better align with real-world clinical decision-making [11]. Clinicians typically begin by ordering lab tests informed by patients’ symptoms and medical history, interpreting the results, and making a diagnostic decision [16]. Inspired by this process, we introduce a lab-informed pretraining strategy, in which lab tests are treated as intermediate clinical signals. It enables the model to perform diagnostic reasoning in a stepwise manner rather than all at once. To enhance the quality of medical knowledge, a diagnosis KG is constructed by selecting highly relevant clinical entities (e.g., diseases, drugs, phenotypes) and incorporating structured triples transformed from unstructured textual information. To mitigate the high computational cost during deployment, we decouple the KG from the model’s prediction phase. This design allows the use of large-scale knowledge graphs without introducing additional inference-time overhead, thereby improving scalability and efficiency.
Therefore, we propose DuaLK, a Dual-Expertise Synergy framework, that collaboratively enhances representation by both Lab-informed pretraining and Knowledge-guided learning for diagnostic reasoning. A diagnosis-specific KG is constructed from multiple clinical databases by using large language models (LLMs) to capture hierarchical and semantic relationships among medical concepts. Next, A lab-informed learning module is designed to utilize individual health data through graph neural networks and a customized pretraining proxy task based on lab test data to rectify patient-level embeddings. In general, these two expertise are collaborated within an encoder-decoder architecture, enabling mutual refinement and enhancing clinical prediction accuracy. The main contributions of this paper are as follows:
  1. 1.
    We construct a diagnosis knowledge graph (KG) with a bi-hierarchical structure upon multiple medical knowledge sources to support predictive modeling. Experimental results on two public EHR datasets demonstrate that the proposed KG significantly enhances diagnostic prediction performance (e.g., F1-score and recall), while maintaining compatibility with various baseline models through a flexible and adaptable embedding strategy.
  2. 2.
    To the best of our knowledge, this is the first work to leverage lab test data for pretraining, modeling stepwise diagnostic reasoning through lab assignment and abnormality detection. Furthermore, supplementary experiments reveal that incorporating laboratory information significantly improves predictive performance, highlighting the crucial role of lab tests in clinical diagnosis.
  3. 3.
    We propose a diagnostic prediction framework explicitly designed to fuse both public and individual medical knowledge. Our approach consistently outperforms state-of-the-art baselines, demonstrating strong generalization and adaptability—even as data volume or diagnostic code complexity increases.
@@ 2 Related Work
Figure 1: Overview of the proposed DuaLK framework.
2.1 Clinical Prediction Models
Deep learning models have been widely applied in predictive healthcare to provide guidance for preventive care. Both RNN-based [17, 18, 19, 20, 21] and CNN-based models [22, 23] have been developed to capture temporal and spatial dependencies for clinical risk predictions. With the rise of graph neural networks (GNNs), many works incorporate ontologies over medical concepts (e.g. diseases) to enrich representations, such as CGL [1], SeqCare [11], and GraphCare [9]. Meanwhile, Transformer-based models like G-BERT [4] and its extensions leverage proxy tasks to extract information from structured EHR data, particularly for patients with limited visit histories [14, 13, 6]. Most of these methods rely on graphs with simple, single-type relations and demonstrate superiority across various EHR prediction tasks, which highlights the potential of enhancing prediction by more complex, multi-relational medical knowledge.
2.2 Knowledge-Guided Clinical Prediction
@@
Predictive models that leverage graph and transformer architectures have achieved state-of-the-art performance in various clinical prediction tasks, highlighting their potential to advance predictive healthcare through the use of medical knowledge. Such knowledge can be categorized into individual knowledge and public knowledge bases, as outlined below:
  • Individual Knowledge: Prior studies such as CGL [1] and Sherbet [5] construct EHR graphs based on interactions (e.g., disease complications) among medical codes in admissions, capturing hidden associations for clinical prediction. Transformer-based models like G-BERT [4] also adopt customized pretraining tasks to exploit extra supervision signals from EHR data, thereby improving convergence and enhancing clinical prediction [24, 6]. However, existing methods either rely on oversimplified associations among medical concepts or employ proxy tasks disconnected from clinical reasoning, limiting their ability to model diagnostic decision processes.
@@
  • Public Knowledge: External knowledge is derived from public data sources beyond the EHR dataset. Researchers have explored simple ontologies such as the ICD-9 hierarchy [10], retrieval-augmented generation (RAG) methods with unstructured medical descriptions like RAM-EHR [12], and large-scale ontologies for prediction, such as in GraphCare [9]. However, they often emphasize either hierarchical or semantic knowledge in isolation, which may introduce noisy or misaligned information and impair prediction performance.
While public and individual knowledge each provide complementary insights, most existing studies rely on only one of them [15]. Effectively combining both hierarchical and semantic knowledge sources is essential for improving the robustness and generalizability of predictive healthcare models.
2.3 Clinical Prediction with Multi-aspect Features
@@
In addition to leveraging medical-domain knowledge, combining both structured and unstructured features also enhance predictions. MiME [25] and GCT [2] use graph structures with lab results to augment representations, while CGL [1] and MedGTX [26] leverage unstructured clinical notes to highlight the value of additional information. However, most approaches rely heavily on these multi-aspect features, limiting their applicability to data without those features. Our proposed framework integrates laboratory features as auxiliary data in graph construction and proxy tasks, enhancing predictions even when no desired feature is available.
2.4 Reasoning in Clinical Prediction
Many studies treat graph learning as a form of reasoning based on neighboring entities [5, 11, 9], which often reduces reasoning to static associations. Reinforcement learning-based methods attempt to emulate clinical inquiry processes through multi-step interactions [27]; however, their reliance on iterative decision-making leads to high training complexity and substantial data requirements, making them difficult to scale and less applicable in real-world clinical settings [28]. More recently, LLMs have been directly used as predictors to simulate clinician-like reasoning [29, 30], a strategy which is prone to hallucinations in medical contexts and involves high computational costs. Overall, most of them fall short in explicitly modeling the stepwise and causally structured nature of clinical reasoning, underscoring the need for more faithful and efficient reasoning frameworks.
@@ 3 Methodology
In this section, we propose DuaLK, a dual-expertise framework designed for predictive healthcare that aims to provide comprehensive diagnoses based on public and individual knowledge. An overview of DuaLK is shown in Figure 1, and the framework operates in three main steps:
Step 1: Build a diagnosis-specific KG enriched by hierarchies upon diseases and phenotypes via LLM, then provide initial embeddings as priors for diseases.
Step 2: Generate a disease-lab EHR graph for graph learning, and customize a lab-informed proxy task to provide stepwise guidance for prediction.
Step 3: Get patient-level embeddings through the encoder-decoder structure, refining public knowledge by lab results for accurate clinical prediction.
@@ 3.1 General Notation
An EHR dataset $\mathcal{S}$ is a collection of temporal admission records of $N$ patients $\{\mathbf{p}_{1},\mathbf{p}_{2},...,\mathbf{p}_{N}\}$, and $\mathbf{p}_{u}$ denotes patient $u$. Such patient can also be represented as a sequence of $T_{u}$ admission records $(V_{1}^{u},V_{2}^{u},..,V_{T_{u}}^{u})\in\mathbf{p}_{u}$ in chronological order. We omit the patient index $u$ in the following sections and explain our framework using a single patient as an example for clarity. We denote the entire set of medical concepts as $\mathcal{C}=\{c_{1},c_{2},...,c_{|\mathcal{C}|}\}$ where $|\mathcal{C}|$ is the vocabulary size of medical concepts. Each admission, like $t$-th admission $V_{t}\in(V_{1},V_{2},..,V_{T})$, contains a subset of $\mathcal{C}$. Note that, We consider diseases and lab results as medical concepts in proposed framework, and we denote diseases $\{d_{1},d_{2},...,d_{|\mathcal{D}|}\}$ and lab results $\{l_{1},l_{2},...,l_{|\mathcal{L}|}\}$ within vocabulary $\mathcal{D},\mathcal{L}\subset\mathcal{C}$.
Figure 2: Overview of Augmented Bi-hierarchical KG.
3.2 Generate Diagnosis KG
We establish knowledge scholar by converting both existing medical KGs and verified textual knowledge into the structured triplet set via the in-context-learning power of LLMs. Figure 2 illustrates how to augment KG with textual information. In general, we aim to integrate and standardize a uniform KG based on both medical ontologies and textual knowledge.
HKG Construction
Healthcare KGs (HKG) integrate information from multiple datasets [31], providing a more complex structure than all-in-one ontologies such as UMLS [32]. We focus on constructing a healthcare KG that links diseases, drugs, and phenotypes to support precise clinical prediction. Therefore, five standardized ontologies and experimentally derived data are adopted for KG construction, ensuring broad coverage across key medical entities:
  1. 1.
    DrugBank [33]: The comprehensive resource containing detailed pharmaceutical knowledge. We retrieved the latest version (5.1.12), published on March 14, 2024. Our focus is on synergistic drug interactions, which represent the bidirectional connections between two drugs.
  2. 2.
    DrugCentral [34]: The verified database that provides drug-disease relations, including indications, contraindications, and off-label uses. We utilized the updated SQL Database, released on November 1, 2023, for our study.
  3. 3.
    HPO [35]: The Human Phenotype Ontology offers details on phenotype abnormalities associated with diseases. We used the most recent updates, from April 19, 2024, focusing on disease-phenotype and phenotype-phenotype relationships, and extracted verified associations with expertly curated annotations.
  4. 4.
    ICD-9-CM [10]: The encoding system is to classify medical conditions for billing purposes, and it underlies the diagnostic information in most EHR data. For our purposes, we extracted the parent-child relationships among codes to describe disease interactions.
  5. 5.
    SIDER [36]: The Side Effect Resource provides side effects caused by various drugs. We retrieved this information from the data release dated October 21, 2015.
The HPO and the ICD-9-CM hierarchies serve as the foundation for our KG, with bi-hierarchical structure enhancing its robustness and connectivity. We denote the resulting healthcare KG, sourced from existing databases, as $\mathcal{G}_{M}$, containing $M$ tuples of head, relation, and tail.
LLM-KG Generation
We further extract relational triples from text information by leveraging the in-context learning capabilities of LLMs. Text information is parsed from the following four clinical websites/databases:
  1. 1.
    Mayo Clinic [37] maintains a knowledge base with clinical information on over 2000 diseases. We collected Mayo Clinic web knowledge and mainly focus on both “Symptoms & Causes” (Overview, Symptoms, When to see a doctor, Causes, Risk Factors, Complications, and Prevention) and “Diagnosis & Treatment” parts for each condition.
  2. 2.
    Orphanet [38] is a clinical database dedicated to rare diseases. We collected data focusing on information about definitions, prevalence, treatment, and clinical descriptions for rare diseases.
  3. 3.
    Rare Disease Database [39] contains supplementary knowledge source for uncovered rare diseases. We collected data to retrieve information on disease overviews, symptoms, causes, treatments, and clinical trials.
  4. 4.
    Wikipedia is the most well-known online encyclopedia. It designs a specific webpage for ICD-9-CM codes [^1] The Wikipedia ICD-9-CM webpage: https://en.wikipedia.org/wiki/List_of_ICD-9_codes. We retrieved related information for each disease on the separate webpage. For example, the page of “Typhoid Fever” provides detailed information on symptoms, causes, diagnosis, prevention, treatment, epidemiology, history, terminology, and societal impact, offering more comprehensive and diverse knowledge compared to other textual sources.
Incorporating unstructured textual information into the Diagnosis KG is achieved by leveraging the in-context learning capabilities of LLMs to extract relational triples from curated disease descriptions. As all inputs originate from verified or expert-curated sources, the risk of hallucination is minimal. A prompting-based approach is adopted instead of retrieval-augmented generation (RAG), showing better empirical performance—likely due to reduced semantic loss and the relatively low task complexity. The LLM-generated triples in the format [ENTITY 1, RELATION, ENTITY 2] are aggregated to form a knowledge graph $\mathcal{G}_{N}$ containing $N$ triples.
To further improve output robustness, two strategies are introduced during inference: (1) Iterated Prompting, where the full input is re-prompted to ensure completeness; and (2) Re-read Prompting, where the same prompt is passed repeatedly for refinement. Final triples are manually reviewed and added to the Diagnosis KG as additional semantic edges, enriching disease node representations. LLM-KG is performed using Llama-3.1-8b [40], an open-source language model selected for its accessibility and compatibility with local deployment. However, other models like DeepSeek [41] are also compatible with the proposed workflow. A detailed prompt example is provided in Appendix5.4.
Diagnosis-KG Integration
To construct a unified Diagnosis KG $\mathcal{G}_{H}$, we integrate two components: the HKG $\mathcal{G}_{M}$ and the LLM-generated KG $\mathcal{G}_{N}$, as defined below: